# Non-IID data in Federated Learning: A Systematic Review with Taxonomy, Metrics, Methods, Frameworks and Future Directions

### v1

@@ Abstract
Recent advances in machine learning have highlighted Federated Learning (FL) as a promising approach that enables multiple distributed users (so-called clients) to collectively train ML models without sharing their private data. While this privacy-preserving method shows potential, it struggles when data across clients is not independent and identically distributed (non-IID) data. The latter remains an unsolved challenge that can result in poorer model performance and slower training times. Despite the significance of non-IID data in FL, there is a lack of consensus among researchers about its classification and quantification. This systematic review aims to fill that gap by providing a detailed taxonomy for non-IID data, partition protocols, and metrics to quantify data heterogeneity. Additionally, we describe popular solutions to address non-IID data and standardized frameworks employed in FL with heterogeneous data. Based on our state-of-the-art review, we present key lessons learned and suggest promising future research directions.
@@ I Introduction
As data becomes increasingly decentralized (a.k.a. federated), residing on edge devices such as smartphones or within isolated institutional silos (so-called clients), the limitations of traditional ML approaches have become more pronounced. Data privacy regulations, such as the General Data Protection Regulation (GDPR) [3] and the Health Insurance Portability and Accountability Act (HIPAA) [4], impose substantial requirements on collecting and processing personal data, making the centralized aggregation of such data impractical or legally infeasible.
@@ I-A Motivation
Typically, for this synthetic data partitioning, centralized datasets such as CIFAR-10 [8], MNIST [9] or FMNIST [10] are divided into non-overlapping subsets, which are then distributed independently to individual clients. Usually, various types of data skewness, such as label imbalance, feature distribution divergence, and quantity skew, can be systematically generated through specific partition protocols. We use the term partition protocol to refer to the systematic partitioning functions used to divide a dataset into smaller subsets, leading to a federated data environment. These partition protocols are crucial in simulating meaningful experimental scenarios for FL. Depending on the protocol, the resulting distributions can either follow an IID (independent and identically distributed) scheme, where each client receives data that mirrors the overall distribution of the dataset, or a non-IID scheme (a.k.a. non-IIDness), where data is distributed in a skewed or heterogeneous manner across clients.
The choice of a partition protocol significantly impacts the performance and convergence of FL algorithms, as non-IID data often presents more significant challenges for model generalization and parameter aggregation than an IID data setting. Thus, designing and selecting appropriate partition protocols is critical to evaluating the robustness and effectiveness of FL methods.
@@ I-B Previous surveys on non-IID data in FL
However, despite the critical importance of establishing a solid understanding of the existing types of data skewness in FL and the necessity of using standardized approaches to simulate realistic data partitions, few studies have thoroughly addressed the need to create a meaningful and comprehensive categorization of existing data partitioning techniques in FL. While we have identified a few surveys that provide foundational insights into the challenges of data heterogeneity, our work advanced beyond these contributions by offering an extended taxonomy alongside a more detailed overview of existing protocols and metrics.
@@ I-C Contribution
This survey represents the first systematic literature review dedicated to organizing and synthesizing the existing knowledge of data distribution skewness in the context of data heterogeneity in FL. The conducted review allows us to present a thorough analysis of the various types of data skewness explored in existing research alongside a detailed description and categorization of state-of-the-art partition protocols. Furthermore, we systematically examine the non-IID metrics employed in FL studies to evaluate the degree of data heterogeneity.
This review not only builds upon previous work, providing a more robust overview of the prevalent methods and metrics for simulating non-IID schemes, but also provides a collection of good practices to promote consistency in experimental setups, enable fair comparison across studies, and ensure that FL models are rigorously tested under conditions that accurately reflect practical deployment environments.
@@ I-D Paper structure
The rest of this survey is organized as follows: Section II provides an overview of the basic notions of FL. Section III describes the systematic literature review approach utilized to compile the results of this survey. Section IV categorizes the various types of data skewness considered in the state of the art. Section V overviews the landscape of partition protocols and heterogeneity metrics utilized in previous research. Section VI reports existing state-of-the-art solutions to tackle non-IIDness in FL. Section VII collects information about existing standardized frameworks for FL and their capacities to deal with heterogeneous data. Section VIII compiles a collection of lessons learned and a list of good practices based on the knowledge acquired during the execution of this work. Section IX summarizes the future work vision from the various research efforts analyzed. Finally, Section X synthesizes the key findings from our systematic literature review and overviews the potential directions for future work in the field. Fig. 1 illustrates the organizational structure of our systematic review.
Figure 1: Outline of our systematic review
@@ II Federated learning basics
FL is an innovative ML paradigm that enables training models on distributed datasets across multiple clients holding local data samples without exchanging them. This approach contrasts traditional centralized ML techniques, where all the data get aggregated in one location for model training.
@@
$\min_{w}l(w):=\sum_{k=1}^{K}{\frac{n_{k}}{n}L_{k}(w)}$ (1)
where $w$ represents the parameters (a.k.a. weights) of the global model trained, $n_{k}$ represents the size of $\mathcal{D}_{k}$, and $n_{k}$ is the total number of examples held by all participating clients. $L_{k}(w)$ is a local objective function such as cross-entropy loss for a supervised classification task [11].
The core idea behind FL is to train the global model $w$ by aggregating the local model updates ((e.g., model gradients or weights) from the participating clients. The latter occurs through a process called Federated Averaging (FedAvg[1, 18], where each device trains its local model on its data, and the central server aggregates the model updates from the devices to update the global model. This iterative process of local training and global model aggregation continues for a pre-defined number of communication rounds ($T$), allowing the model to be trained on a diverse range of data sources while preserving the privacy of the individual data contributors, as the raw data never leaves the local devices.
@@ II-A2 Cross-devices
Cross-device FL operates on a much larger scale, involving numerous edge devices such as smartphones, IoT devices, or wearables. This type of FL deals with many participants with limited computational resources and smaller, often intermittently available datasets. Cross-device FL must handle challenges like unreliable network connections, device heterogeneity, and frequent client unavailability. Its use is commonly in mobile keyboard prediction, voice assistants, and other consumer-facing applications [20].
@@ II-A5 Vertical FL
Vertical FL, or attribute-based FL, is used when participants have datasets with identical sample IDs but different feature spaces. This scenario often arises in multi-party business collaborations. For example, a bank and an e-commerce company might have different data types (attributes or features) for the same set of customers [24].
@@ II-B Non-IID data in FL
In centralized learning, non-IID data typically refers to variations within a single dataset that violate the assumption of IID [17, 25]. The latter might include class imbalances, temporal shifts in data distribution, or biases in data collection. Although challenging, researchers can often address such issues using data augmentation [26], resampling [27], or careful stratification during train-test splits [28]. In contrast, non-IID data in FL presents a more complex scenario. In FL, not only can each participant’s local dataset be non-IID, but the data distribution across participants can also be highly heterogeneous. It leads to a two-level non-IID problem: intra-client non-IID (within each participant’s dataset) and inter-client non-IID (across different participants). The decentralized nature of FL, combined with privacy constraints that limit direct data sharing, makes traditional centralized approaches to handling non-IID data insufficient [29].
@@ II-C Downstream effects of non-IID data in FL
The resulting challenges and complexities make non-IID data in FL attractive to researchers. In this context, Fig.2 depicts how research focusing specifically on non-IID challenges in FL has notably increased since 2020, also peaking in 2023, highlighting the growing attention this issue is receiving. It is essential to emphasize that the apparent decrease in 2024 is due to the timing of our data collection, as papers were gathered during the third quarter, leading to incomplete data for that year (see Section III)). Thus, although general FL research continues to expand more quickly, the consistent rise in studies addressing non-IID data shows that it is emerging as a critical topic within the FL research community [40].
III Literature review methodology
In this section, we outline the methodology used to identify and select the relevant papers and articles that form the basis of our literature review. Additionally, we present a detailed meta-analysis of these papers, highlighting compelling information and insights.
III-A Systematic review
We employed a systematic review (SR) based on the PRISMA methodology [41] to conduct a thorough and structured analysis of the current literature on non-IID-ness in FL. We improved the latter methodology by considering top-tier published papers and research from high-level universities. The SR approach involved gathering, filtering, and assessing relevant academic papers and resources from various scholarly databases. The review process began with an identification phase, as illustrated in Fig. 3. We utilized six reputable academic repositories: Google Scholar [42], IEEE Xplore [43], PubMed [44], Scopus [45], and Web Of Science [46]. We chose these platforms for their comprehensive coverage across multiple academic fields and their extensive collection of research publications. Our search strategy involved 83 queries across ten categories related to non-IID-ness in FL, including general non-IID-ness, label skew, attributes skew, quantity skew, distribution shift, non-IID metrics, performance metrics, partitioning protocols, datasets, and frameworks. We input each of the 83 queries into each academic repository to retrieve relevant papers. After eliminating duplicate documents, we compiled a collection of 5489 unique papers for further analysis.
@@ III-A Systematic review
We systematically screened each article after the initial collection to identify potentially valuable papers. Our screening process consisted of the following steps for each paper retrieved:
@@
The latter allowed us to systematically filter the large pool of papers, focusing on those most likely to contribute to our study of non-IID-ness in FL.
@@
Some relevant papers are often not published in conferences or journals. Still, it is worth analyzing whether they have received many citations or come from high-ranking universities. Then, 33 additional papers are marked as ”useful” if it was not published in a journal or conference (i.e., published in Arxiv), but any of the following occurs:
@@
III-B Meta-analysis of retrieved literature
We comprehensively examined the 235 papers selected through our SR process to provide a general glance at the current research on non-IID-ness in FL. Through this analysis, we seek to contextualize individual findings within the broader landscape of non-IID FL literature and provide a foundation for understanding the collective progress and challenges in this rapidly evolving domain.
@@ IV Taxonomy of data heterogeneity in FL
Figure 7: Taxonomy for non-IIDness in FL
@@
Writing $\mathcal{D}_{i}\sim\mathbb{D}_{i}$ for the $i$th clients’ local data, and denoting the probability density function (pdf) of $\mathcal{D}_{i}$ by $f^{(i)}$, data skew means that
@@
In the supervised learning setting with $\mathcal{D}_{i}=(y_{i},x_{i})$, where $y_{i}$ are the labels and $x_{i}$ the attributes, we can further divide general data skew into label and attribute skews as in the following:
@@
    i.e., the marginal or conditional distributions of the labels differ between some pairs of clients. In general, with heterogeneous data we might find either marginal or conditional skew or both. However, if there is no attribute skew (see below), then marginal label skew implies conditional label skew and vice versa. More formally, assuming $f_{X}^{(i)}=f_{X}^{(j)}$ and $f_{X|Y}^{(i)}=f_{X|Y}^{(j)}$ (no attribute skew) implies $f_{Y}^{(i)}\not=f_{Y}^{(j)}\Leftrightarrow f_{Y|X}^{(i)}\not=f_{Y|X}^{(j)}$. The proof follows immediately from Bayes’ Theorem:
@@
Unlike previous surveys and systematic reviews, we include the notion of Modality skew under the taxonomy of non-IIDness for the first time. It is a specific type of data skew referring to a setting where different clients have data from varying input modalities, such as text, images, audio, tabular, or sensor data, as opposed to all clients having data of the same type. This occurs when clients collect or generate data in distinct formats, leading to heterogeneity in the input data types across the network. Modality skew is a subtype of the attribute and label skews whose interdependence is detailed in Fig. 8.
@@
    As a result of participation skew, not all clients contribute equally or consistently to model training. In the presence of data skew, such imbalances can give rise to biases in the global model that reflect the over-representation of certain clients while others are underutilized or excluded. This can lead to various issues, such as different forms of unfairness (for example, when some group of clients has systematically worse utility) and problems with out-of-distribution (OOD) generalization (e.g., when there is data skew between training and testing/production set of clients).

### v2

# Non-IID data in Federated Learning: A Survey with Taxonomy, Metrics, Methods, Frameworks and Future Directions

@@ Abstract
Recent advances in machine learning have highlighted Federated Learning (FL) as a promising approach that enables multiple distributed users (so-called clients) to collectively train ML models without sharing their private data. While this privacy-preserving method shows potential, it struggles when data across clients is not independent and identically distributed (non-IID) data. The latter remains an unsolved challenge that can result in poorer model performance and slower training times. Despite the significance of non-IID data in FL, there is a lack of consensus among researchers about its classification and quantification. This technical survey aims to fill that gap by providing a detailed taxonomy for non-IID data, partition protocols, and metrics to quantify data heterogeneity. Additionally, we describe popular solutions to address non-IID data and standardized frameworks employed in FL with heterogeneous data. Based on our state-of-the-art survey, we present key lessons learned and suggest promising future research directions.
@@ I Introduction
As data becomes increasingly decentralized (a.k.a. federated), residing on edge devices such as smartphones or within isolated institutional silos (so-called clients), limitations of traditional ML approaches such as privacy and efficiency have become more pronounced. Data privacy regulations, such as the General Data Protection Regulation (GDPR) [3] and the Health Insurance Portability and Accountability Act (HIPAA) [4], impose substantial requirements on collecting and processing personal data, making the centralized aggregation of such data impractical or legally infeasible.
@@
Throughout this paper, we highlight key findings from our survey using light blue background text to facilitate quick identification of significant insights.
@@ I-A Motivation
Typically, for this synthetic data partitioning, centralized datasets such as CIFAR-10 [8], MNIST [9] or FMNIST [10] are divided into non-overlapping datasets, which are then assigned to individual clients. Usually, various types of data skewness, such as label imbalance, feature distribution divergence, and quantity skew, can be systematically generated through specific partition protocols. We use the term partition protocol (a.k.a partition method) to refer to the systematic partitioning functions used to divide a dataset into smaller subsets, leading to a federated data environment. These partition protocols are crucial in simulating meaningful experimental scenarios for FL. Depending on the protocol, the resulting distributions can either follow an IID (independent and identically distributed) scheme, where each client receives data that mirrors the overall distribution of the dataset, or a non-IID scheme (a.k.a. non-IIDness), where data is distributed in a skewed or heterogeneous manner across clients.
The choice of a partition protocol significantly impacts the performance and convergence of FL algorithms, as non-IID data often presents additional challenges for model generalization and parameter aggregation compared to an IID data setting. Thus, designing and selecting appropriate partition protocols is critical to evaluating the robustness and effectiveness of FL methods.
@@ I-B Previous surveys on non-IID data in FL
However, despite the critical importance of establishing a solid understanding of the existing types of data skewness in FL and the necessity of using standardized approaches to simulate realistic data partitions, few studies have thoroughly addressed the need to create a meaningful and comprehensive categorization of existing data partitioning techniques in FL. While we have identified a few surveys that provide foundational insights into the challenges of data heterogeneity, our work advances beyond these contributions by offering an extended taxonomy alongside a more detailed overview of existing protocols and metrics.
@@ I-C Contribution
This work is the first technical survey dedicated to organizing and synthesizing the existing knowledge regarding distribution skewness in FL, providing a comprehensive taxonomy and detailed categorization of data heterogeneity types. We specifically contribute by:
  1. 1.
    Introducing a taxonomy that organizes and synthesizes the diverse forms of non-IIDness.
  2. 2.
    Reviewing state-of-the-art partition protocols for simulating non-IID scenarios in FL.
  3. 3.
    Methodically analyzing non-IID metrics for quantifying data heterogeneity, fostering standardized evaluations.
  4. 4.
    Offering good practices to ensure experimental consistency, fair benchmarking, and realistic deployment testing.
Our work bridges gaps in existing surveys, serving as a critical reference for researchers and practitioners in FL.
@@ I-D Paper structure
The rest of this survey is organized as follows: Section II provides an overview of the basic notions of FL. Section III describes the approach utilized to retrieve the most relevant papers this survey. Section IV categorizes the various types of data skewness considered in the state of the art. Section V overviews the landscape of partition protocols and heterogeneity metrics utilized in previous research. Section VI reports existing state-of-the-art solutions to tackle non-IIDness in FL. Section VII collects information about existing standardized frameworks for FL and their capacities to deal with heterogeneous data. Section VIII compiles a collection of lessons learned and a list of good practices based on the knowledge acquired during the execution of this work. Section IX summarizes the future work vision from the various research efforts analyzed. Finally, Section X synthesizes the key findings from our survey and overviews the potential directions for future work in the field. Fig. 1 illustrates the organizational structure of our survey.
Figure 1: Outline of our survey
@@ II Federated learning basics
FL is an innovative ML paradigm that enables training models on distributed datasets across multiple clients holding local data samples (a.k.a examples, individuals, data points) without exchanging them. This approach contrasts traditional centralized ML techniques, where all the data get aggregated in one location for model training.
@@
$\min_{w}l(w):=h(L_{k}(w))$ (1)
where $w$ represents the parameters (a.k.a. weights) of the global model trained, and $h$ is function to aggregate the clients parameters. $L_{k}(w)$ is a local objective function such as cross-entropy loss for a supervised classification task [11].
The core idea behind FL is to train the global model $w$ by aggregating the local model updates (e.g., model gradients or weights) from the participating clients. The latter occurs through a process called Federated Averaging (FedAvg[1, 18], where each device trains its local model on its data, and the central server aggregates the model updates from the devices to update the global model. In FedAvg the aggregation function $h(L_{k}(w))=\sum_{k=1}^{K}{\frac{n_{k}}{n}L_{k}(w)}$, where $w$ represents the parameters of the global model trained, $n_{k}$ represents the size of $\mathcal{D}_{k}$, and $n$ is the total number of samples held by all participating clients. The mentioned iterative process of local training and global model aggregation continues for a pre-defined number of communication rounds ($T$), allowing the model to be trained on a diverse range of data sources while preserving the privacy of the individual data contributors, as the raw data never leaves the local devices. Alternative aggregation functions to FedAvg in FL are explained in Section VI.
@@ II-A2 Cross-devices
Cross-device FL operates on a much larger scale, involving numerous edge devices such as smartphones, IoT devices, or wearables. This type of FL deals with many participants with limited computational resources and smaller, often intermittently available datasets. Cross-device FL must handle challenges like unreliable network connections, device heterogeneity, and frequent client unavailability. Its use is commonly in mobile keyboard prediction, voice assistants, and other consumer-facing applications [20, 21].
@@ II-A5 Vertical FL
Vertical FL, or attribute-based FL, is used when participants have datasets with identical samples but different feature spaces. This scenario often arises in multi-party business collaborations. For example, a bank and an e-commerce company might have different data types (attributes or features) for the same set of customers [25].
II-A6 Peer-to-peer FL
Also known as decentralized FL, removes the central server, using peer-to-peer communication for model aggregation via epidemic or diffusion processes [26]. This approach improves scalability, fault tolerance, and privacy, making it ideal for edge networks or blockchain-based systems. Unlike hierarchical FL, it relies on flexible, pairwise, or group-based updates.
@@ II-B Non-IID data in FL
In centralized learning, non-IID data typically refers to variations within a single dataset that violate the assumption of IID [17, 27]. The latter might include class imbalances, temporal shifts in data distribution, or biases in data collection. Although challenging, researchers often address such issues using data augmentation [28], resampling [29], or careful stratification during train-test splits [30]. In contrast, non-IID data in FL presents a more complex scenario. In FL, not only each participant’s local dataset can be non-IID, but the data distribution across participants can also be highly heterogeneous. It leads to a two-level non-IID problem: intra-client non-IID (within each participant’s dataset) and inter-client non-IID (across different participants). The decentralized nature of FL, combined with privacy constraints that limit direct data sharing, makes traditional centralized approaches to handling non-IID data insufficient [31].
@@ II-C Downstream effects of non-IID data in FL
  • Susceptibility to attacks: Non-IID data can aggravate vulnerabilities to cybersecurity threats unique to FL, such as poisoning attacks, where adversarial participants introduce malicious updates to compromise the global model [39].
@@
The resulting challenges and complexities make non-IID data in FL relevant for researchers. In this context, Fig.2 depicts how research focusing specifically on non-IID challenges in FL has notably increased since 2020, also peaking in 2023, highlighting the growing attention this issue is receiving. It is essential to emphasize that the apparent decrease in 2024 is due to the timing of our data collection, as papers were gathered during the third quarter, leading to incomplete data for that year (see Section III)). Thus, although general FL research continues to expand more quickly, the consistent rise in studies addressing non-IID data shows that it is emerging as a critical topic within the FL research community [43].
III Literature identification and meta-analysis
In this section, we outline the methodology used to identify and select the relevant papers and articles that form the basis of our survey. Additionally, we present a detailed meta-analysis of these papers, highlighting compelling information and insights.
III-A Methodology for identifying relevant literature
We employed a methodology for identifying relevant literature based on the PRISMA methodology [44] to conduct a thorough and structured analysis of the current literature on non-IID-ness in FL. We improved the latter methodology by considering top-tier published papers and research from high-level universities. The approach involved gathering, filtering, and assessing relevant academic papers and resources from various scholarly databases. The process began with an identification phase, as illustrated in Fig. 3. We utilized six reputable academic repositories: Google Scholar [45], IEEE Xplore [46], PubMed [47], Scopus [48], and Web Of Science [49]. We chose these platforms for their comprehensive coverage across multiple academic fields and their extensive collection of research publications. Our search strategy involved 83 queries across ten categories related to non-IID-ness in FL, including general non-IID-ness, label skew, attributes skew, quantity skew, distribution shift, non-IID metrics, performance metrics, partitioning protocols, datasets, and frameworks. We input each of the 83 queries into each academic repository to retrieve relevant papers. After eliminating duplicate documents, we compiled a collection of 5489 unique papers for further analysis.
@@ III-A Methodology for identifying relevant literature
We screened each article after the initial collection to identify potentially valuable papers. Our screening process consisted of the following steps for each paper retrieved:
@@
The latter allowed us to filter the large pool of papers, focusing on those most likely to contribute to our study of non-IID-ness in FL.
@@
Some relevant papers are often not published in conferences or journals. Still, it is worth analyzing whether they have received many citations or come from high-ranking universities. Then, 36 additional papers are marked as ”useful” if it was not published in a journal or conference (i.e., published in Arxiv) and any of the following occurs:
@@
III-B Meta-analysis of relevant literature
We comprehensively examined the 235 papers selected through our process to provide a general glance at the current research on non-IID-ness in FL. Through this analysis, we seek to contextualize individual findings within the broader landscape of non-IID FL literature and provide a foundation for understanding the collective progress and challenges in this rapidly evolving domain.
@@ IV Taxonomy of data heterogeneity in FL
Figure 7: Taxonomy for non-IIDness in FL. The figure categorizes different types of data skews and heterogeneities. Percentages indicate the proportion of each type across global, cross-silo-related, and cross-device-related papers.
@@
Writing $\mathcal{D}_{i}\sim\mathbb{D}_{i}$ for the $i$th clients’ local data, where $\mathbb{D}_{i}$ represents the underlying data distribution from which the $i$th client’s local data $\mathcal{D}_{i}$ is sampled, and denoting the probability density function (pdf) of $\mathcal{D}_{i}$ by $f^{(i)}$, data skew means that
@@
In the supervised learning setting with $\mathcal{D}_{i}=(y_{i},x_{i})$, where $y_{i}$ are the labels (target outputs or classes assigned to samples) and $x_{i}$ the attributes (features or properties describing samples), we can further divide general data skew into label and attribute skews as in the following:
@@
    i.e., the marginal $f_{X}^{(i)}$ or conditional $f_{Y|X}^{(i)}$ distributions of the labels differ between some pairs of clients. In general, with heterogeneous data we might find either marginal or conditional skew or both. However, if there is no attribute skew (see below), then marginal label skew implies conditional label skew and vice versa. More formally, assuming $f_{X}^{(i)}=f_{X}^{(j)}$ and $f_{X|Y}^{(i)}=f_{X|Y}^{(j)}$ (no attribute skew) implies $f_{Y}^{(i)}\not=f_{Y}^{(j)}\Leftrightarrow f_{Y|X}^{(i)}\not=f_{Y|X}^{(j)}$. The proof follows immediately from Bayes’ Theorem:
@@
Unlike most previous surveys, we include the notion of Modality skew under the taxonomy of non-IIDness. It is a specific type of data skew referring to a setting where different clients have data from varying input modalities, such as text, images, audio, tabular, or sensor data, as opposed to all clients having data of the same type. This occurs when clients collect or generate data in distinct formats, leading to heterogeneity in the input data types across the network. Modality skew is a subtype of the attribute and label skews whose interdependence is detailed in Fig. 8.
@@
    As a result of participation skew, not all clients contribute equally or consistently to model training. In the presence of data skew, such imbalances can give rise to biases in the global model that reflect the over-representation of certain clients while others are underutilized or excluded. This can lead to various issues, such as different forms of biases (for example, when some group of clients has systematically worse utility) and problems with out-of-distribution (OOD) generalization (e.g., when there is data skew between training and testing/production set of clients).